Makenzie2021:

The probability of choosing a green grape out of the bag that contains red and green grapes without looking is StartFraction 11 over 15 EndFraction. Which term best describes this probability? impossible unlikely likely certain

jhonyy9:

the probability is \[\frac{ 11 }{ 15 }\] of choosing a green grape out of the bag what result from this fraction ?

jhonyy9:

@supie any idea ?

supie:

Well, a good way to do this is to get the precent out of 100 so we would have a percentage chance of the probability - just a way I like to do it we need to get the denominator to 100 and 15 x 6.6 = 99, which is close enough to 100, so, \(\frac{11}{15}\times\frac{6.6}{6.6}\) multiply straight across 15 x 6.6 = 100 11 x 6.6 72.6 / 73 So we get \(\frac{11}{15}\times\frac{6.6}{6.6}=\frac{72.6}{100}\) so there is a 72.6% chance We should automatically know what `impossible` is not a valid option because if it was impossible then there would be a zero precent chance, 0/100 so we can eliminate `impossible` if it was unlikely then it would be somewhere under 50/50%/50/100, what we have now is about 73/73%/73/100, so I think we can say that `unlikely` can also be eliminated from our answer options. Do you agree, @jhonyy9?